OpenStudy (anonymous):

what are the factors of the following trinomial ? 2x^2 -7x+ 3

myininaya (myininaya):

ax^2+bx+c find two factors of a*c that multiply to be a*c add up to be b so we have a*c=2*3=-6(-1) and -6-1=-7 we replace -7x with -6x-1x so we have 2x^2-6x-1x+3 now factor by grouping 2x(x-3)-1(x-3) (x-3)(2x-1)

OpenStudy (anonymous):

soo the answer ( 2x-1) ( x+3)

myininaya (myininaya):

no (x-3)(2x-1) or since multiplication is commutative (2x-1)(x-3)